Numerical Analysis I, Fall 2009 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 10

Assigned Nov 20, 2009. Due Nov 27, 2009.

1. Section 5.1: Problems 2(d,e,f), 4, 10, 11.

For problem 11: Ignore the sentence ‘Are these results still valid ...’ 2. Section 5.2: Problems 5(a), 6(a), 12, 13, 14, 15, 16, 19, 21.

3. Section 5.3: Problems 1, 4, 6, 12.

For Problem 1, also report I − I_{3} and I − I_{4}.